In the following reaction, 5.96 g of Ammonia (NH3) reacts completely with 180.00g Carbon dioxide (CO2). Which is the limiting reagent?

Given: RMM of N=14; RMM of  H=1; RMM of C=12; RMM of O=16. 

                                       2 NH + CO→ CN2OH4 + H2O     

[L]


1
Expert's answer
2022-02-14T17:30:04-0500

"RMM(NH_3)=14+3\times1=17"

"RMM(CO_2)=12+2\times16=44"


"n(NH_3)=5.96\ g\times\frac{1\ mol}{17\ g}=0.35\ mol"

"n(CO_2)=180.00\ g\times\frac{1\ mol}{44\ g}=4.1\ mol"


According to the equation, NH3 and CO2 react in 2 : 1 molar ratio. The actual ratio is 0.35 : 4.1, which is less than 2 : 1. Therefore,


NH3 is the limiting reactant.